# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v3

@@ Abstract
Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality. Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond a few state variables ($\sim$3-6 limit in prior studies). To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework. PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to directly inform neural network-based policy learning. A key contribution is our highly efficient Projected PG-DPO (P-PGDPO) variant. This approach uniquely utilizes BPTT to obtain rapidly stabilizing estimates of the Pontryagin costates and their crucial derivatives with respect to the state variables. These estimates are then analytically projected onto the manifold of optimal controls dictated by PMP’s first-order conditions, significantly reducing training overhead and enhancing accuracy. This enables a breakthrough in scalability: numerical experiments demonstrate that P-PGDPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables). Critically, the framework accurately captures complex intertemporal hedging demands, a feat often elusive for other methods in high-dimensional settings. P-PGDPO delivers near-optimal policies, offering a practical and powerful alternative for a broad class of high-dimensional continuous-time control problems.
@@ 1 Introduction
High-dimensional optimal control problems often fall prey to the “curse of dimensionality” (Bellman, , 1966), whereby the computational cost of classical dynamic programming (DP) approaches grows exponentially with the number of state variables. In a continuous-time finance setting, this difficulty becomes starkly evident in multi-asset Merton problems (Merton, , 1973), where an investor dynamically allocates wealth among numerous risky assets whose drift and volatility may depend on exogenous state variables such as macroeconomic indicators or stochastic volatility factors. While DP or Hamilton–Jacobi–Bellman (HJB) PDE solvers provide a rigorous framework for low-dimensional cases (Fleming and Soner, , 2006; Bertsekas and Tsitsiklis, , 1995), they become intractable as the number of assets or state dimensions increases. This difficulty is pervasive, impacting not only continuous-time settings (which are often addressed via discretization for numerical solution) but also analogous discrete-time models where similar state-space explosion issues arise. Consequently, prior influential studies in these domains are typically limited to analyzing problems involving up to approximately 3-5 state variables (Campbell and Viceira, , 1999; Lynch, , 2001; Balduzzi and Lynch, , 1999) or specific frameworks handling around 6 state variables (Garlappi and Skoulakis, , 2010), often necessitating simulation-based approaches for tractability (Brandt et al., , 2005).
A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong and Zhou, , 2012). Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint). Notably, adding one more state dimension does not produce an exponential explosion in computational requirements, in contrast to DP (Bellman, , 1966; Campbell et al., , 2003). This makes Pontryagin-based methods appealing for large-scale portfolio optimization problems involving many assets and exogenous state processes (Karatzas et al., , 1987).
Recent advances in deep learning further enhance the scalability of methods for solving high-dimensional stochastic control problems (e.g., Han et al., , 2018; Raissi et al., , 2019, for general PDE-based approaches). Concurrently, deep learning techniques are being applied directly to financial dynamic programming and stochastic control. Some approaches tackle infinite horizon problems via the HJB equation (Duarte et al., , 2024), or solve the HJB equation for utility maximisation under constraints using formulations based on second-order backward stochastic differential equations (2BSDEs) (Davey and Zheng, , 2022). Others aim to directly solve associated nonlinear PDEs using methods like deep branching algorithms (Nguwi et al., , 2024). While these BSDE-based methods and direct PDE solvers are powerful, applications such as the Merton HJB have revealed limitations in accuracy or efficiency for certain variants (Nguwi et al., , 2024). Moreover, some works effectively leverage the Stochastic Maximum Principle (SMP) and adjoint BSDEs, often through duality methods, to handle even non-Markovian problems in finance (Davey and Zheng, , 2022). Alongside these PDE-centric or BSDE-centric deep learning approaches, another line of research, more akin to reinforcement learning, focuses on directly parameterizing the control policy with neural networks and optimizing the expected objective function through simulation and gradient ascent (e.g., Zhang and Zhou, , 2019; Kopeliovich and Pokojovy, , 2024, for portfolio problems with transaction costs or under Heston dynamics). Our PMP-based PG-DPO framework builds upon similar ideas of direct policy optimization but is distinctively guided by the continuous-time optimality conditions of PMP. It utilizes BPTT-derived costate estimates for finite-horizon problems, aiming for enhanced scalability and a novel way to enforce optimality conditions for problems with sufficient regularity.
One notable strategy related to the use of adjoint methods in optimal control, and central to our work, is to directly parameterize the policy (i.e., consumption and investment controls) with neural networks, then apply backpropagation-through-time (BPTT) to enforce Pontryagin’s optimality conditions. This strategy, which we term the baseline “Pontryagin-Guided Direct Policy Optimization (PG-DPO),” circumvents the need to approximate value functions or PDE solutions. Instead, it uses simulation-based gradients that automatically incorporate the adjoint processes, drastically reducing computational overhead in high-dimensional settings (Kushner and Yin, , 2003; Borkar and Borkar, , 2008).
The purpose of this paper is to develop and implement this PG-DPO framework for the general multi-asset, state-dependent continuous-time portfolio problem. We consider an investor who must decide on a continuous consumption rate and an allocation to $n$ risky assets with state-dependent drift and volatility. By parameterizing the policy with neural networks, we can capture potentially complex relationships dependent on both wealth and the exogenous state. We show how a single forward pass in simulation, coupled with a backward pass via automatic differentiation (BPTT), yields unbiased estimates of the costate processes that PMP requires, guiding the policy towards optimality without the drawbacks of DP. This approach fundamentally relies on the existence and accurate estimation of the costate processes dictated by PMP, applicable whenever the control problem satisfies appropriate regularity conditions allowing for a well-defined Hamiltonian system.
Furthermore, we introduce the “Projected PG-DPO” (P-PGDPO) variant, which significantly enhances efficiency and accuracy. This variant exploits the observation that BPTT-derived costate estimates often stabilize much faster than the full policy network. Once the costates are sufficiently stable (after a short warm-up phase), one can bypass the potentially suboptimal neural policy outputs. Instead, these costate estimates are directly substituted into the PMP first-order conditions (FOCs), effectively projecting the learned costate information onto the manifold of controls satisfying PMP’s necessary conditions. This recovers a near-optimal control analytically at every time-state point with minimal additional training, significantly reducing computational burden. While other deep learning methods such as Deep BSDE are also applied to high-dimensional problems, their ability to accurately capture complex, state-dependent optimal policy structures, particularly the crucial intertemporal hedging components, has been limited in challenging financial applications. Our numerical experiments (Section 4) demonstrate that P-PGDPO not only scales to problems with dozens of assets and state variables (e.g., n=50, k=10 in our experiments) but, critically, it is capable of precisely disentangling and reproducing these vital hedging demands—a feat not typically achieved by existing approaches in such high-dimensional settings. To the best of our knowledge, this explicit and accurate recovery of complex hedging terms in a high-dimensional, PMP-guided direct policy optimization framework is a novel contribution. This superior performance in both scalability and fidelity to the true optimal policy, especially its nuanced components, highlights the transformative potential of leveraging the PMP structure via BPTT-derived costate estimates for tackling a broad class of complex, high-dimensional control problems where amenability to costate estimation under suitable regularity conditions exists.
@@
2 Multi-Asset Continuous-Time Portfolio Problem with State Variables
In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1973; Kim and Omberg, , 1996; Liu, , 2007), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously. Each risky asset features its own drift and volatility, leading to an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}$. Furthermore, an exogenous state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ influences market parameters such as returns and interest rates, adding additional complexity to the investor’s decision.
Although the wealth $X_{t}$ remains scalar, the investor’s optimal policy now depends on both $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and a high-dimensional state $\mathbf{Y}_{t}$. We handle this through Pontryagin’s maximum principle (Pontryagin, , 2014; Yong and Zhou, , 2012), introducing a costate for $X_{t}$ and another for $\mathbf{Y}_{t}$. Numerically, one can represent $\{\bm{\pi}_{t},C_{t}\}$ via neural networks, thereby enabling policy gradient or dynamic programming algorithms even in multi-asset, state-dependent environments.

### v4

@@ Abstract
Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality. Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond a few state variables ($\sim$3-6 limit in prior studies). To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework. Our framework accurately captures both myopic demand and complex intertemporal hedging demands, a feat often elusive for other methods in high-dimensional settings. P-PGDPO delivers near-optimal policies, offering a practical and powerful alternative for a broad class of high-dimensional continuous-time control problems. PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to directly inform neural network-based policy learning. A key contribution is our highly efficient Projected PG-DPO (P-PGDPO) variant. This approach uniquely utilizes BPTT to obtain rapidly stabilizing estimates of the Pontryagin costates and their crucial derivatives with respect to the state variables. These estimates are then analytically projected onto the manifold of optimal controls dictated by PMP’s first-order conditions, significantly reducing training overhead and enhancing accuracy. This enables a breakthrough in scalability: numerical experiments demonstrate that P-PGDPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables).
@@ 1 Introduction
The theory of dynamic consumption and portfolio choice has been a cornerstone of financial economics, providing fundamental insights into asset pricing and long-term investment behavior (Merton, , 1973; Campbell and Viceira, , 1999, 2001). Despite its theoretical elegance and contributions to our understanding, the theory faces a major obstacle in practical implementation: the lack of scalable methods for solving high-dimensional portfolio choice problems involving many assets and state variables.
The Bellman equation associated with such problems is subject to the curse of dimensionality, and obtaining solutions in realistic settings is often infeasible. Existing solution methods are typically complex, ad hoc, and do not scale well with dimensionality. This is one reason why Markowitz’s static mean-variance framework remains dominant in practice, despite its limitations and the greater realism offered by dynamic models. [^1] Cochrane wrote “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice. ” (Cochrane, , 2022, p.4)
In this paper, we make a technological contribution by proposing a solution approach that leverages machine learning with neural networks to solve high-dimensional dynamic portfolio choice problems, where both the number of assets $n$ and the number of state variables $k$ are large. There are two primary challenges inherent in such problems. The first is the well-known curse of dimensionality: computational methods often fail to scale efficiently as the dimensionality of the problem increases in $n$ and $k$. The second, and more subtle, challenge arises from the specific structure of dynamic portfolio choice itself. Unlike other applications of reinforcement learning—such as Atari, Go, or autonomous driving—where achieving a near-optimal policy is often sufficient for strong empirical performance, portfolio optimization requires much greater precision. A well-established theoretical benchmark exists, dating back to the seminal work of Merton, (1973), which demonstrates that the optimal portfolio can be decomposed into two critical components: the myopic demand and the intertemporal hedging demand. The myopic component represents the demand for a mean-variance efficient portfolio given current investment opportunities, while the intertemporal component hedges against expected future changes in those opportunities driven by evolving economic conditions. Accurately identifying both components is essential for evaluating investor behavior and designing long-term investment strategies—yet it remains a formidable challenge, particularly in high-dimensional settings. Our proposed approach is specifically designed to address both challenges: overcoming the curse of dimensionality and enabling the accurate decomposition of optimal portfolios into their myopic and intertemporal components.
We adopt a continuous-time framework in which asset return means and covariances are driven by a $k$-dimensional state vector, with the state variables evolving as Itô diffusions, following Merton, (1973). This modeling approach benefits from the rich toolkit of stochastic calculus—tools generally unavailable in discrete-time formulations. [^2] See Duarte et al., (2024) for a discussion of the advantages of continuous-time models over discrete-time models. In the continuous-time setting, (quasi-)analytic solutions are available under affine dynamics, where the Hamilton–Jacobi–Bellman (HJB) equation reduces to a system of ordinary differential equations (Kim and Omberg, , 1996; Liu, , 2007; Buraschi et al., , 2010). However, for general non-affine dynamics, no universal solution method exists. To address this limitation, we introduce a machine learning–based approach that applies to arbitrary market dynamics. Neural networks, known for their ability to approximate complex high-dimensional functions, are used to parameterize the policy directly, enabling us to capture intricate dependencies on both wealth and the exogenous state variables—even when the dimension $k$ is large.
While we focus on Brownian motion as the source of return risk, the methods can be extended to accommodate jumps. Furthermore, if a discrete-time model is viewed as an approximation of a continuous-time model, our approach can be adapted accordingly.
Our approach leverages Pontryagin’s Maximum Principle (PMP) as its theoretical foundation, in contrast to the more commonly used dynamic programming (DP) framework based on the Hamilton–Jacobi–Bellman (HJB) equation. Building on PMP, we propose two methods: Pontryagin-Guided Direct Policy Optimization (PGDPO) and its variant, Projected PGDPO (P-PGDPO).
PMP characterizes optimal controls through a coupled system of forward and backward differential equations. The forward equation governs the evolution of wealth and state variables driving asset returns, which we implement via Monte Carlo simulation. This allows for direct computation of the expected utility and avoids the need to estimate the value function across future periods—a core but computationally intensive requirement of DP-based approaches.
The backward equation governs the dynamics of the costates (or adjoint variables), which correspond to the first-order partial derivatives of the value function with respect to the state variables. Modern machine learning frameworks such as PyTorch allow us to compute these costates and their derivatives, the second-order partial derivatives of the value function, automatically through backpropagation through time (BPTT) and automatic differentiation. We demonstrate that a single forward simulation pass, coupled with BPTT, yields unbiased estimates of the costate processes required by PMP. As a result, our approach avoids the need for explicit solution of the backward equation or approximation of the value function, thereby sidestepping major bottlenecks of DP. Costates—as well as the value function—emerge as natural byproducts of the learning process.
In PGDPO, the control policy—comprising consumption and portfolio allocations—is parameterized by neural networks. The expected utility, defined as the discounted sum of utility from consumption and terminal wealth, is then directly maximized using stochastic gradient ascent, where gradients are computed using BPTT. A key computational advantage of our approach lies in its complete avoidance of HJB residual minimization, where the residual is defined as the cumulative squared error of the HJB equation. Computing the HJB residual typically requires evaluating first- and second-order partial derivatives of the value function at each time–state grid point and aggregating the errors over time, incurring high computational cost. Duarte et al., (2024) mitigate this by replacing full gradient computation with the second derivative of a carefully constructed univariate function. In contrast, our method bypasses HJB residual computation entirely, resulting in significant computational savings.
Conceptually, our approach shares similarities with recent methods that parameterize control policies via neural networks and directly optimize the expected objective function through simulation and gradient ascent, thereby circumventing the need to solve the HJB equation (Zhang and Zhou, , 2019; Kopeliovich and Pokojovy, , 2024, e.g., for portfolio problems with transaction costs or Heston dynamics). However, our PMP-based PGDPO framework distinguishes itself by being explicitly guided by continuous-time optimality conditions derived from PMP.
Moreover, unlike standard reinforcement learning methods such as Proximal Policy Optimization (PPO) or Trust Region Policy Optimization (TRPO), which rely on sampled estimates of policy gradients, our method computes exact gradients via BPTT. This yields lower-variance updates and enhances computational efficiency, particularly in high-dimensional settings (Kushner and Yin, , 2003; Borkar and Borkar, , 2008).
The second method, P-PGDPO, builds on PGDPO by substituting the estimated derivatives of the value function directly into the structural first-order conditions from PMP. These conditions take the form derived originally by Merton, (1973), explicitly decomposing the optimal portfolio into myopic and intertemporal hedging demands. We demonstrate the numerical accuracy of our methods by comparing the solutions they produce to a quasi-analytic benchmark under affine dynamics, where the true solution can be obtained by solving a system of ordinary differential equations.
The two components of the optimal portfolio—myopic and intertemporal hedging demands—are not merely of theoretical interest; their accurate identification requires high-fidelity solutions. As shown by Constantinides, (1986), transaction costs generate only second-order utility losses, allowing portfolios to deviate substantially from the frictionless optimum while still appearing nearly optimal in terms of utility. Similarly, Cochrane, (1988) demonstrates that first-order deviations from the optimal policy can result in only second-order losses in utility.
This disconnect between utility value and policy structure poses a major challenge for traditional reinforcement learning algorithms. These methods often converge to policies that are near-optimal in utility but structurally far from the true optimum—particularly in high-dimensional settings. In such cases, neural networks must search for optimal policies over an $n$-dimensional control space, where $n$ denotes the number of risky assets. The resulting exponential growth in complexity renders standard policy optimization techniques computationally infeasible as $n$ increases. Our numerical experiments confirm that increasing the number of risky assets $n$ presents a significantly greater challenge than increasing the number of state variables $k$ in high-dimensional settings.
The P-PGDPO method is motivated by the empirical observation that value function derivatives—i.e., costates—are typically easier to learn than the policy functions themselves. In practice, these partial derivatives tend to stabilize much earlier during training. Once the costate estimates have stabilized—usually after a short warm-up phase—the method can bypass the potentially suboptimal outputs of the policy networks. Instead, the costates and their derivatives are directly substituted into the first-order conditions (FOCs) from PMP, whose structural form closely mirrors the conditions originally derived by Merton, (1973). This effectively projects the learned costate information onto the manifold of controls that satisfy PMP’s necessary conditions. The resulting control policy adheres closely to the structural form of the true optimal solution dictated by PMP. It is recovered quasi-analytically at each time–state point, requiring minimal additional training and incurring substantially reduced computational cost. Our PMP-based PGDPO framework thus leverages BPTT-derived estimates of costates and their derivatives, enhancing scalability while providing a principled mechanism for enforcing first-order optimality conditions.
We provide a theoretical justification for the P-PGDPO method in Theorem 1. When the objective function surface is relatively flat near the optimum, the first-order condition (FOC) violation for the policy obtained in the warm-up stage tends to be small. The theorem states that if this FOC deviation is small and the costate estimates obtained via BPTT are sufficiently accurate, then the control policy recovered in the second stage is close to the true optimal policy.
We demonstrate the accuracy of our numerical methods by solving a problem in which the state process follows a multi-factor Ornstein–Uhlenbeck (OU) process—a special case of affine dynamics—for which quasi-analytic benchmark solutions can be obtained by solving a system of ordinary differential equations (see Section 2.3.2). We also compare our methods with the Deep Backward Stochastic Differential Equation (Deep BSDE) approach, which is representative of existing techniques for solving high-dimensional partial differential equations, particularly the Hamilton–Jacobi–Bellman (HJB) equations arising in DP. In addition, we apply our second method, P-PGDPO, to a non-affine setting by adding non-linear terms to the expected returns of risky assets, thereby demonstrating its applicability to a broader class of dynamics.
Our numerical experiments (Section 4) show that our first method, PGDPO, produces solutions comparable to those of the Deep BSDE method in terms of mean-squared errors. In contrast, the second method, P-PGDPO, yields significantly smaller errors. This improvement arises because PGDPO inherits the common limitation of policy optimization algorithms—difficulty in discovering the true optimum—whereas P-PGDPO overcomes this issue by directly exploiting the structural form of the first-order conditions from Pontryagin’s Maximum Principle (PMP).
P-PGDPO not only scales to problems involving dozens of assets and state variables (e.g., $n=50$, $k=10$), but also accurately disentangles and reconstructs complex intertemporal hedging demands. Furthermore, our results demonstrate that the method generates stable solutions in non-affine settings and converges smoothly to the benchmark affine solution as the coefficients of the nonlinear terms approach zero.
By contrast, the Deep BSDE approach, while successful in other high-dimensional applications such as option pricing and hedging, fails to capture the full structure of the optimal policy in our setting—particularly the intertemporal hedging components.
The method’s superior performance—both in scalability and in fidelity to the full structure of the optimal policy—highlights the transformative potential of leveraging PMP through BPTT-derived costate estimates for a broad class of high-dimensional control problems that admit stable costate representations.
There is a substantial literature on dynamic portfolio choice, aimed at obtaining concrete solutions either analytically or numerically. With the exception of the continuous-time models with affine dynamics discussed above, most existing contributions are formulated in discrete time. For example, Campbell and Viceira, (1999) and Campbell and Viceira, (2001) employ log-linear approximations to the dynamic budget constraint to obtain approximate closed-form solutions. Balduzzi and Lynch, (1999), Lynch and Balduzzi, (2000), and Lynch, (2001) apply discrete-space approximations to solve portfolio choice problems with transaction costs. Garlappi and Skoulakis, (2010) develop a DP approach based on state variable decomposition.
Brandt et al., (2005) propose a Monte Carlo simulation method for a discrete-time portfollio choice model that projects simulated values of relevant functions—such as the marginal value of wealth—onto a pre-specified set of basis functions. In this sense, their method shares similarities with the forward simulation and auto-differentiation steps in our PG-DPO framework. However, their approach does not explicitly leverage PMP or utilize the representational power of neural networks. As a result, it is not easily scalable and lacks the flexibility for generalization. In contrast, our method is PMP-guided, scalable, and extensible to broader classes of control problems.
In continuous time, there has been significant technological advancement in solving partial differential equations, which can be applied to the HJB equation arising from DP. Notable methods include Deep BSDE, Deep Galerkin, Physics-Informed Neural Networks (PINNs), Deep Splitting, and Deep Branching, among others (Han et al., , 2018; Sirignano and Spiliopoulos, , 2018; Raissi et al., , 2019; Beck et al., , 2021; Nguwi et al., , 2024).
In parallel, substantial progress has been made in solving optimal control problems through the development of policy evaluation, policy gradient, $q$-learning, and PPO/TRPO methods in stochastic environments (Wang et al., , 2020; Jia and Zhou, 2022a, ; Jia and Zhou, 2022b, ; Jia and Zhou, , 2023; Zhao et al., , 2023). Wang and Zhou, (2020); Dai et al., (2023); Huang et al., (2024) apply these techniques to dynamic mean-variance analysis, while Dai et al., (2025) use them to solve a Merton-type portfolio optimization problem. These approaches involve learning the asset return process from observed return data. Geng et al., (2023) propose the Factor Learning Portfolio Optimization (FaLPO) method for solving a Merton-type portfolio optimization problem with stochastic factors. Their approach combines policy learning with model calibration. By contrast, we assume that the data-generating process is known and do not address the problem of learning from data. Instead, our focus is on a more general consumption and portfolio choice problem involving many risky assets and multiple state variables, whereas their analysis is limited to a small number of risky assets and state variables.
Davey and Zheng, (2022) develop a BSDE-based method for portfolio optimization under portfolio constraints that is also based on the maximum principle, similar in spirit to our work. However, their method requires minimizing the terminal value error (as in Deep BSDE), maximizing the Hamiltonian at each state and time, and solving a BSDE system. In contrast, our method directly maximizes expected utility without penalizing approximation errors and without solving the backward equation explicitly, resulting in a significantly simpler and more efficient approach.
More recently, Duarte et al., (2024) propose a machine learning approach for infinite-horizon continuous-time problems using DP, while Cheridito et al., (2025) develop a method that combines generalized policy iteration with PINNs to solve finite-horizon continuous-time DP problems. By contrast, our method is based on PMP, rather than DP. The aforementioned DP-based methods avoid the explicit evaluation of first- and second-order partial derivatives of the value function by instead computing the second derivative of a univariate function when assessing the HJB residual. In contrast, our approach computes these partial derivatives directly—via automatic differentiation of costate variables—providing crucial information for accurately capturing intertemporal hedging components. This ability to compute the value function’s derivatives with high accuracy enables precise estimation of both the intertemporal hedging terms and the myopic component of the optimal portfolio. [^3] Cochrane emphasizes the critical role of partial derivatives in determining the optimal portfolio for long-term investors, particularly when discussing the implementation challenges of Merton’s dynamic portfolio theory—a point we partially quoted earlier: “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice… What are the partial derivatives of the investor’s value function that guide state-variable investments? That is harder still…” (Cochrane, , 2022, pp. 4–5). To our knowledge, the two aforementioned papers do not demonstrate accurate recovery of intertemporal hedging terms in settings with both large numbers of assets and state variables.
To our knowledge, the existing literature has demonstrated the ability to solve problems involving at most a dozen risky assets and a dozen state variables—with two notable exceptions. Davey and Zheng, (2022) consider a setting with 20 assets, but under a constant investment opportunity set, i.e., without state variables influencing asset returns. [^4] Our companion paper demonstrates that our method is applicable to constrained portfolio optimization problems involving up to 1000 assets (Huh et al., , 2025). Cheridito et al., (2025) address a case with 25 assets, though without providing evidence regarding the accuracy of the resulting solution. In contrast, our method substantially improves scalability, handling much larger values of $n$ and comparably high values of $k$, while preserving accuracy even in high-dimensional settings.
@@
2 Multi-Asset Continuous-Time Portfolio Choice Problem with State Variables
We consider the continuous-time portfolio choice problem introduced by Merton, (1973), in which a representative investor allocates wealth across $n$ risky assets and a risk-free asset, while consuming continuously over time. The investor selects an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}=(\pi_{1,t},\ldots,\pi_{n,t})^{\top}$, where $\pi_{i,t}$ denotes the proportion of current wealth $X_{t}$ invested in the $i$-th risky asset at time $t$ and notation denotes the transpose of a vector or a matrix.
The investment opportunities are characterized by an exogenous $k$-dimensional state process $\mathbf{Y}_{t}=(Y_{1,t},\ldots,Y_{k,t})^{\top}\in\mathbb{R}^{k}$, which governs the drift and volatility of asset returns. For instance, $\mathbf{Y}_{t}$ could include macroeconomic indices, volatility factors, exchange rates, etc. The investor’s optimization problem involves dynamically choosing both the portfolio weight vector $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and the consumption rate $C_{t}$, conditional on the current state $\mathbf{Y}_{t}$ and wealth $X_{t}$.
To analyze this problem, we apply Pontryagin’s maximum principle (Pontryagin, , 2014; Yong and Zhou, , 2012), introducing costate variables corresponding to both the wealth process $X_{t}$ and the state process $\mathbf{Y}_{t}$. From a computational perspective, the control variables $\{\bm{\pi}_{t},C_{t}\}$ can be parameterized using neural networks, facilitating the implementation of policy gradient methods or DP algorithms in high-dimensional, state-dependent settings.